[Halld-jlab] Hall D Staff Meeting Notes, January 14, 2013

Mark M. Ito marki at jlab.org
Tue Jan 15 15:20:09 EST 2013


Find them at 
https://halldweb1.jlab.org/wiki/index.php/Private:Hall_D_Staff_Meeting,_January_14,_2013 
and below.
   __

Private:Hall D Staff Meeting, January 14, 2013

    Present: Eugene Chudakov (chair), Hovanes Egiyan, Mark Ito, David
    Lawrence, Bert Manzlak, Mike McCaughan, Lubomir Pentchev, Yi Qiang,
    Dmitry Romanov, Elton Smith, Sasha Somov, Simon Taylor, Ivan
    Tolstukhin, Beni Zihlmann

Announcements, Eugene

      * some items in the project plan are late: online, BCAL, FDC
           + costs about 20% over
           + 1 M$ contingency will be needed
           + 7 M$ per year over cost on average
           + directors reviews on trends in Jan. and Feb.
      * will do an ETC to analyze what needs to be spent to finish project
           + note that ETC change request is for a single fiscal year
           + new one a full analysis to end of project
           + DOE rejected Hall C change request, process stuck for now
      * jobs for us:
           + estimate how much we need for completion
           + DOE requires that we present basis of estimate
           + use justification for old values or new values with new
             justification?
      * end of April: DOE Lehman review
      * after Lehman review, re-baseline
           + old schedule dumped, start from scratch with new schedule

Engineering, Eugene

      * magnet at 80 K
      * leak in 4th coil, nitrogen leak, probably in the chimney, liter of
        ice per year
      * problem with water cooling at CHL, burst water pipe underground,
        could be weeks before repair, we are not affected directly

FDC, Lubomir

      * end of last week finished package #[?]
      * three cells installed last week at Blue Crab on package #[?]
           + installed three more this week
           + wire broken in cleaning, but fixed now
           + biggest distance from wire to strong-back is 50 microns
      * 4th package:
           + done before, moved to EEL last week, flushed with gas, HV
             applied on Friday, no problems
           + Beni started testing today
      * Dave Butler + 2 more technicians back out at Blue Crab
           + now working on cathodes
      * Beni: 1st test results, all wires working, problem with TDC in
        set-up
           + 2 running FADC-125 with old firmware, will read out each plane
             in three steps to check if all strips are working

Safety, Bert

      * review last week for solenoid power up
           + recommendations given to engineering group
           + document for cool-down, electronics, and cryogenics: operators
             need to read and sign
      * test set-ups outside hall, e. g., EEL, storage building, TEDF, Blue
        Crab, all need work control document
           + refer from Hall D list

Calorimetry, Elton

FCAL

      * frame delivered last December and adjusted
      * John Frye from IU, technical engineer, arrived this morning with
        additional pieces
      * hope to start stacking tomorrow, 2 students are here, Dan and
        Manual
      * Matt will be here Wed. afternoon for a day
      * IU team coordinating stacking effort with Tom
      * work will be done on downstream platform

BCAL

      * will start gluing process soon, getting organized now
      * need to review production process
      * today collecting materials
           + light guides (50% on site)
           + BCAL monitoring boards
                o 20 strings, enough for 2 modules
           + Tedlar light guide covers being made by hand now
                o plan to stamp them out, die in hand, Jim Fochtman will
                  get job done off-site
                o need to be formed around light guide
      * re-did the wand
           + screws to each light guide
           + pick up 10 light guides, place close to face of BCAL, apply
             glue, tighten down one-by-one
           + there is a plate on the BCAL with mounting screws
      * read-out packages
           + all mechanical pieces ordered, many came in in December
           + some have to be anodized, procurement for that contract held
             up
           + hardware to put it together has been ordered but has not
             arrived yet, all should be here by end of month
           + pre-assembly required before assembly
                o bend cooling pipes, put connectors on, glue ends to
                  gluing plate, glue RTV[?] boards onto cooling plant
                o get manpower for this in Feb., Tom will arrange
                o put things together in March, Fernando and Nick
                o assembly will take couple of months
                o boards arrived, populated, setting jumpers now, Chris
                     # 6 mV was biggest outlier, 10 mV tolerance
      * will put light guides on and use mini-BCAL set-up, run cosmics
        overnight
           + then put modules back in box and move them to Hall

Beamline, Sascha

      * checking properties for fixed array PMTs and active divider
           + before holiday break, note posted reporting measurements
           + tube looks good, timing resolution 200 ps
           + high rates tested, linear up to 10 MHz
           + got quotes from Hamamatsu, reasonable, $440 each, no base
           + so far so good
      * Vladimir Popov developing amplifier
           + first version tested on Friday
           + some oscillations seen, fixed
           + if successful, may consider active base for TOF
           + active divider works OK though amplitude is small and may need
             amplification
           + Vladimir says components are very cheap
      * UConn contract awarded, readiness review scheduled for Feb. 7
      * microscope: Hovanes has idea for reading out voltages and current
        from groups of SiPMs
           + will bring Fernando to next beamline meeting
      * contract with Catholic: Franz received RFP, some legal objections

Pair Spectrometer, Ivan

      * new fixture to allow access for cleaning surface on side of
        scintillator
           + Anatoly has altered design
           + problem was that fixture, if pressed too tight, squeezes glue
             out

Tagger Magnet, Eugene

      * Magnet supposed to arrive next month
      * Bill will visit the company making the vacuum chamber
           + contractor is late, hopefully arrive in Feb.

Accelerator, Mike

      * ion chamber discussions underway
           + putting two ion chambers around goniometer, also one around
             the beam dump
      * active collimator interface meeting was held
      * idea to add harps to the system
           + already some in front of radiator
           + add downstream of tagger and in front of detector[?]

Online, Hovanes

      * sometime in August, specs sent to Richard for EPICS interface to
        UConn control system
           + Richard has sent responses
           + no monitoring of voltage or current for SiPM bias in current
             plan, discussion on how to add this if at all
      * counting house
           + had meeting today, Dave L., Dave A., Diane Napier, Jason
             Willoughby, Dave Fazenbaker
           + Dave F. presented sketch
           + only three seats in sketch, will go for 5
           + price estimates will be obtained
      * VME boards
           + buying 7 CPUs, 4 quad core, 3 dual core
           + 57 others will be ordered later
           + 11 on-hand already
           + Glenn and David Abbott running order, waiting for I7 modules
      * online computers
           + four computers installed, Dave L., Dave A., Elliott studying
           + getting experience with infiniband
           + Dave L. working on a level 3 software prototype
      * meeting Wed. with Computer Center on computing for counting house,
        will transmit specifications to them
           + we know about infiniband, level 3 farm nodes and event
             builders (EMU) 75% of online computers
           + will need some for web server, MySQL server
           + after CPUs evaluated, will look at RAID needs

Offline

Tracking, Simon

      * geometry changes, added extra material
      * bench-marking of code, running valgrind
      * looking at data challenge files

BCAL reconstruction, David

      * BCAL reconstruction on hold, waiting on Regina to complete work on
        sampling fluctuations
           + Andrei promises some results soon

Recent Offline Issues, Mark

      * CCDB roll-out underway
      * data challenge 1 done
           + working out system for distributing data, will rely on grid
             tools
      * re-configuring offline tasks
           + need to revisit our task list
           + old task list good for getting the software started,
             detector-by-detector emphasis
           + now need to focus on specific reconstruction areas that don't
             necessarily break down detector-by-detector

    Retrieved from
 
"https://halldweb1.jlab.org/wiki/index.php/Private:Hall_D_Staff_Meeting,_January_14,_2013"

-- 
Mark M. Ito
Jefferson Lab
marki at jlab.org
(757)269-5295


More information about the Halld-jlab mailing list